OpenStudy (anonymous):

(2+3/(x+1))/(1/x+x+x^2 ) in simplifing this complex fraction is the answer ((2x+5)x)/((x+1)1+x^3+x^2 )

OpenStudy (anonymous):

I didn't get that :p. try this- when dividing fractions, multiply by the reciprocal ex. 1/2 div 2/3 = 1/2 times 3/2

OpenStudy (anonymous):

(2+3/(x+1))/(1/x+x+x^2 ) = (2x+2+3)/((x+1)/x+x+x^2 ) [Multiply top and bottom by (x+1) =(2x+5)/((x+1)/x+x+x^2 ) [Simplify Numerator] =x(2x+5)/(1 + x +x^2+x^3) [Multiply top and bottom by x] same?

OpenStudy (anonymous):

this is a complex fraction and i am simplifiying only it should read as this in the attached file

OpenStudy (anonymous):

so the answer should be

OpenStudy (anonymous):

I couldn't read the file. but ((2x+5)x)/((x+1)1+x^3+x^2 ) is not completely simplified, the denominator can lose the parenth around (x+1) and the one multiplying that term is redundant: =x(2x+5)/(1 + x +x^2+x^3)